OpenStudy (anonymous):

Part A: Solve C = (F - 32) for F. Part B: Determine the value of F when C = 8 °C. Part C: Solve -np - 80 > 60 for n. Show your work.

OpenStudy (radar):

I Believe that Part A Formula for C is incomplete.

OpenStudy (anonymous):

c means celsius . its correct.

OpenStudy (radar):

The point is, the formula as is, is incomplete, there is another relationship needed (other than the 32 degree.

OpenStudy (anonymous):

ohhh i see. Opps!! there was a fraction before the parentheses. 5/9

OpenStudy (radar):

Now you're cooking. C = 5/9 (F-32) we want to solve for F. First I would multiply both sides by 9/5 We would then have (9/5) (C) = F - 32. Now add 32 to both sides getting F =(9/5) (C) + 32. Do you agree?

OpenStudy (radar):

When you understand how we converted the equation to solve for F, we will start Part B.

OpenStudy (anonymous):

would we multiply 5/9 by c or use distributive property?

OpenStudy (anonymous):

i was taught to use distributive but you may have another way I can learn.

OpenStudy (radar):

Earth to katlove14, you still there?

OpenStudy (anonymous):

i wrote back. is it not showing up?

OpenStudy (radar):

I see it now, computer problems! I do not want to change the way you were taught, distribute the 5/9 getting C=5/9 F - 5/9 (32) Add 5/9 (32) to both sides, getting C + 5/9 (32) = 5/9 F Now multiply both sides by 9/5 F = 9/5 C + 32 Note we come up with same thing.

OpenStudy (anonymous):

agreed!

OpenStudy (radar):

Are ready for B?

OpenStudy (anonymous):

yep! i get it!

OpenStudy (radar):

F = (9/5)(8) + 32 What did you get?

OpenStudy (anonymous):

i got F=47 degrees

OpenStudy (radar):

I got 46.4 degrees F I think we are close enough

OpenStudy (anonymous):

i do to

OpenStudy (radar):

Is Part C connected with Part A and Part B, or is it a new problem?

OpenStudy (anonymous):

I think its a new problem.

OpenStudy (radar):

Yes, -np - 80 > 60 Just copying it, is that right

OpenStudy (anonymous):

yes.

OpenStudy (radar):

I would first add 80 to both sides of inequality sign getting: -np > 140 agreed ?

OpenStudy (anonymous):

yep!

OpenStudy (radar):

Now the tricky part, to isolate n we would need to divide both sides by -p, but since we are dividing by a minus sign, we must reverse the sense of the inequality sign; n < 140/-p agreed ??

OpenStudy (anonymous):

yes. however, the equation did say >60 for n. did they give us n?

OpenStudy (radar):

No they did not give a value for n or p, but they asked for an inequality for n. Obviously if we are given a value for p then n must be less 140 divided by -p lets say p was 70 then n would have to be less than -2, n could be -3 down to - inifinity.

OpenStudy (anonymous):

ok got it.

OpenStudy (radar):

When working with inequality, if you multiply or divide with a negative you must reverse the sense (direction) of the inequality sign, < becomes > ETC,

OpenStudy (anonymous):

yep

OpenStudy (radar):

good luck with your studies.

OpenStudy (anonymous):

thanks for all your help!